the length of a rectangular lot is 5 m longer than the width. The amount of fencing required is 54 m . Find the dimensions of the lot.: the length of a rectangular lot is 5 m longer than the width. The amount of fencing required is 54 m . Find the dimensions of the lot.
Answer by KnightOwlTutor(216)

We are assigning x=width
x+5 =Length
We know that the perimeter=54
we know the formula for calculating the perimeter:w+L+w+L
combining like terms P=2W+2L
P=2X+2(x+5)=54 m
use the distributive property
2x+2x+10
=54m
Combine like terms
4x+10m=54m
subtract 10 from both sides
4x=44
Divide both sides by 4
x=11
the width =11m
Length=11+5=16m
Double check your answer by plugging in values into the perimeter equation
2(11)+2(16)=
22+32=54=54